# Quantum Monte Carlo simulations in the restricted Hilbert space of Rydberg atom arrays

### v4

@@ 3.2 Local segment update
As the acceptance rate is not necessarily an indicator of good sampling for the observables of interest, we study the correlation in simulation time of the dimer density (rod coverage discussed above) with and without this update, in the presence of all other updates. We choose this observable as inserting and deleting rods directly changes $n_{i}$, which none of the other updates do explicitly. We choose $T=0.15$ and $\Omega=0.6$, as Figs. 3(b), 4(b) and  6(b) show that the acceptance probability for all three off-diagonal updates which we have discussed is close to maximum here, and study the total dimer density at a fixed imaginary time slice ($n=\frac{1}{N}\sum_{i}n_{i}$) starting from various initial conditions. The latter are generated using simulations which employ all updates and are thus representative of the equilibrium distribution. We deliberately do not use parallel tempering here as we want to follow the loss of memory of the initial condition using just the updates available for individual copies of the system. In Fig. 3(c) we show that without the local update, the dimer density does not equilibrate to a unique value as a function of Monte Carlo time for three different initial conditions when averaged over many realizations of the Markov chain dynamics. This implies that the simulation is not ergodic. However, the addition of the local update leads to the a single equilibrium value ,as seen in Fig. 3(c), for the same initial conditions, with a relatively short Monte Carlo time duration required to lose memory of the initial condition. The data presented here is for $L=4$, and the same behavior is expected to hold at higher $L$. This clearly demonstrates the importance of this simple update.
@@ 3.4 Rod diffusion update
Figure 6: (a) A single move within the rod diffusion update shown by projecting it onto the spatial plane. Circled are the numbers of rods at each spatial site which have a temporal overlap with the rod being moved (shown as gray dimer). This move can proceed as proposed as the number of legal forward options ($=1$), is the same as the number of legal backward options ($=1$). (b) The success probability of the rod diffusion update for an $L=8$ lattice. (c) Spatial variance of the dimer density (Eq. (6)) as a function of Monte Carlo time with and without the rod diffusion update for $L=4,T=0.2$ and $\Omega=0.6$.
@@
As explained for the previously described updates, we estimate the region of space-time visited by a single update on average without regarding whether the update is successful or not, and use this estimate to calculate the number of repetitions required to fully sample space-time. Now we can estimate the fraction of space-time which is successfully updated in one Monte Carlo step by our update, and use this to understand the efficiency of the update. We show numerical results for the same in Fig. 6(b) for a range of temperature $T$ and $\Omega/\delta$, and we find that indeed in the two limits discussed above and for values of temperature $\ll\delta,\Omega$, our update has an $O(1)$ success probability. As shown in Fig. 6(b), the inefficiencies begin to be visible only at low temperatures and intermediate values of $\Omega$.
@@
with $\bar{n}=\frac{1}{N}\sum_{i}\braket{n_{i}}_{av}$, where $av$ denotes averaging over a large number of Markov chains. This is calculated as a function of Monte Carlo time. To analyze the efficiency of this update, we choose $\Omega=0.6$ and $T=0.2$, as this is choice of parameters belongs to a regime where the rod diffusion update has a high relative probability (Fig. 6(b)), and the other updates are not efficient (seen in acceptance probabilities in Figs. 3(b),4(b)). Our results are presented in Fig. 6(c), and we see that the local updates are not able to lose memory of the initial condition as the variance does not decay significantly, whereas using the rod diffusion update the variance drops to zero rapidly. The small residual value of the variance is due to the finite number of Markov chains which we have used (as $\braket{n_{i}}_{av}=\bar{n}$ only in the limit of averaging over an infinite number of Markov chains).
@@ 3.6 Benchmarking and autocorrelation functions
To ensure that we have implemented the detailed balance conditions correctly, we perform a comparison with exact diagonalization (ED) for a small system size. For the 24-site periodic lattice shown in Fig. 2, the number of dimer occupation states allowed by the Rydberg constraint are $2649$. This number is small enough to perform a full diagonalization, and thus to calculate energy at finite temperature. We carry out a comparison of the energy for a range of $\Omega$ as shown in Fig. 8(a), and find that the agreement between ED and QMC is of the order of $0.01\%$, and within the error bars of the QMC simulation. Thus we can conclude that the QMC simulations are unbiased and accurate. Note that in Fig. 8(a), we have used $T=0.006$, which is $\ll\Omega$ for the range of $\Omega$ studied.
@@ 4 Phase diagram of Kagomé link lattice
We now study the Kagomé link lattice using ED for the 24-site periodic lattice to gain an understanding of the phase diagram and study larger system sizes using the QMC updates discussed in the previous section. The latter allow us to extract the phase diagram in the thermodynamic limit. The Rydberg atom array on this lattice has been discussed in Ref. [1], where it was found that in the limit of $\Omega\gg\delta$, the ground state is a trivial paramagnet in the language of spin-1/2 particles, and that by lowering $\Omega$, one can cross into a spin liquid phase and eventually into a valence bond solid phase with a large unit cell. These conclusions were reached by carrying out DMRG simulations on quasi-1D cylindrical geometries, and we would like to study the temperature dependence using our QMC directly in the 2D limit. The effect of temperature is most easily understood in the classical limit, where at high temperature, the system accesses all Rydberg allowed configurations, and at $\beta\gg\delta$, the system only accesses the manifold of states which have maximal density. This set is still macroscopically degenerate due to the inherent frustration of the underlying lattice. Once a quantum fluctuation is added, we expect a second energy scale controlled by both $\Omega$ and $\delta$, where the residual classical entropy is released, leading to a ground state with at most $O(1)$ degeneracy.
@@
Using ED, we study the specific heat as a function of both $\Omega$ and $\beta$ (for convenience we set $\delta=1$). We can trace the lower release in entropy as a peak in the specific heat, and we show this in Fig. 9(a). We find that this peak moves smoothly from $T\approx 10^{-5}$ at $\Omega=0.3$ to temperatures of $O(1)$ at high $\Omega$. This suggests that the curve traced by the location of the peak in Fig. 9(a) ends at $T=0$ for $\Omega=0$, which would be the natural expectation. At $\Omega\approx 1$, the distinction between the two peaks in specific heat which signify the release in entropy due to classical constraints and quantum fluctuations is lost. This is expected for $\Omega\approx\delta$ as the separation between scales required for two distinct peaks is lost. This phase diagram suggests that the quantum paramagnet is the correct ground state for all $\Omega>0$. This is in contrast with the expectation of a spin liquid phase. We find that the low lying spectrum for small $\Omega$ has a unique ground state with a roughly equally spaced spectrum above it. This is also inconsistent with a $Z_{2}$ spin liquid on a torus, where one would expect four low lying states which have a topological nature, and are well separated from the rest of the spectrum.
@@
We show our results in Fig. 9(b) and (c), and find that these are similar to the $24$-site lattice for system sizes of $4$ and $8$ down to temperatures of $\approx 0.01$. For low temperatures, calculating $C_{v}$ using the fluctuation- dissipation relationship, i.e. $C_{v}=\beta^{2}(\braket{E^{2}}-\braket{E}^{2})$, has large error bars, and thus we obtain $C_{v}$ directly by taking a numerical derivative of the energy. For our simulation ranges, this ensures an error bar of $<10\%$. The smallest value of $\Omega$ at which this temperature is low enough to detect the second peak in specific heat is $0.7$, and the temperature considered is 30 times smaller than the smallest scale in the system ($\Omega/2=0.35$). This point is consistent with the location of the transition to the spin liquid phase in Ref. [1]. If the spin liquid phase indeed appears close to this transition, it is not seen down to temperatures much smaller than $\Omega,\delta$.
@@ Appendix C Details of rod diffusion update
Each spatial site on the Kagomé link lattice can be visualized as a dimer between two neighboring sites on the Kagomé lattice, and it can be moved by pivoting it about one of its ends, as shown in Fig. 5(a). By extending this into the temporal direction, this pivot corresponds to a movement of a rod. The spatial site of the Kagomé lattice about which this pivot is carried out is chosen to not be in the membrane of modification already developed, as this would lead to the update undoing its own moves. This is similar to the spatial worm update [35], where dimers are pivoted around an exit site, which is not the site via which the worm entered the dimer. For the initial rod, this site is chosen at random, as there is no pre-existing membrane. An example of the membrane of modification is shown in Fig. 5(b). For the Kagomé lattice, a pivoting dimer has three possible options for its new position, as shown in Fig. 5(a). To decide which of these options to choose, we consider the corresponding rod configuration. We calculate the number of conflicting rods generated by moving the current rod to each of these three spatial positions while keeping the temporal ends of the rod the same. We choose the position with the least number of conflicting rods, and denote this number by $n_{c}$. If there are multiple such positions, we store the number of them by denoting them as the number of forward options ($p_{f}$), and choose one at random with uniform probability. If $n_{c}>1$, detailed balance would require a reverse proposed move which coordinates the movement of $n_{c}$ rods into the relevant positions, leading to the reverse movement of the current rod. This would imply that we would have to engineer an update which branches, and carries movements of multiple rods at the same step of the update. This would significantly complicate the update, and we cannot visualize a simple way to do this currently. An example of this sort of movement is shown in Fig. 5(c) and (d), where the movement of the current rod (shown in red) to its new position (shown in blue) creates a conflict with two rods. As shown, these two rods can be on the same spatial site or on different ones.
@@
Thus if $n_{c}>1$, we abort the update. If not, we consider the reverse move, where the rod is being moved from its future position to its current position. In the dimer language, this is now the reverse pivot. Following the same procedure to pick the new position of the rod leads to a calculation of the number of backward options $p_{b}$. Detailed balance now requires that the probability of the forward process, $1/p_{f}$, be equal to that of the reverse process, $1/p_{b}$. This trivially translates to requiring $p_{f}=p_{b}$. If this condition is not met, we abort our update. We relax this condition only for the initial and final rod movement, where we require that $p_{f}^{start}=p_{b}^{end}$ and $p_{b}^{start}=p_{f}^{end}$. A sample pivot process which satisfies detailed balance is shown in Fig. 6(a). If the update is accepted, we update the relevant operators. With each rod movement, we carry all operators ($H_{K}$ and $H_{V}$ type) which are living on it to the next position inhabited by the rod. The update terminates when the current rod is in a position where there are no more conflicting rods to move. A short pseudo-code for this update is presented below.

### v5

@@ 3.2 Local segment update
As the acceptance rate is not necessarily an indicator of good sampling for the observables of interest, we study the correlation in simulation time of the dimer density (rod coverage discussed above) with and without this update, in the presence of all other updates. We choose this observable as inserting and deleting rods directly changes $n_{i}$, which none of the other updates do explicitly. We choose $T=0.15$ and $\Omega=0.6$, as Figs. 3(b), 4(b) and  6(a) show that the acceptance probability for all three off-diagonal updates which we have discussed is close to maximum here, and study the total dimer density at a fixed imaginary time slice ($n=\frac{1}{N}\sum_{i}n_{i}$) starting from various initial conditions. The latter are generated using simulations which employ all updates and are thus representative of the equilibrium distribution. We deliberately do not use parallel tempering here as we want to follow the loss of memory of the initial condition using just the updates available for individual copies of the system. In Fig. 3(c) we show that without the local update, the dimer density does not equilibrate to a unique value as a function of Monte Carlo time for three different initial conditions when averaged over many realizations of the Markov chain dynamics. This implies that the simulation is not ergodic. However, the addition of the local update leads to the a single equilibrium value ,as seen in Fig. 3(c), for the same initial conditions, with a relatively short Monte Carlo time duration required to lose memory of the initial condition. The data presented here is for $L=4$, and the same behavior is expected to hold at higher $L$. This clearly demonstrates the importance of this simple update.
@@ 3.4 Rod diffusion update
Figure 6: (a) The success probability of the rod diffusion update for an $L=8$ lattice. (b) Spatial variance of the dimer density (Eq. (6)) as a function of Monte Carlo time with and without the rod diffusion update for $L=4,T=0.2$ and $\Omega=0.6$.
@@
As explained for the previously described updates, we estimate the region of space-time visited by a single update on average without regarding whether the update is successful or not, and use this estimate to calculate the number of repetitions required to fully sample space-time. Now we can estimate the fraction of space-time which is successfully updated in one Monte Carlo step by our update, and use this to understand the efficiency of the update. We show numerical results for the same in Fig. 6(a) for a range of temperature $T$ and $\Omega/\delta$, and we find that indeed in the two limits discussed above and for values of temperature $\ll\delta,\Omega$, our update has an $O(1)$ success probability. As shown in Fig. 6(a), the inefficiencies begin to be visible only at low temperatures and intermediate values of $\Omega$.
@@
with $\bar{n}=\frac{1}{N}\sum_{i}\braket{n_{i}}_{av}$, where $av$ denotes averaging over a large number of Markov chains. This is calculated as a function of Monte Carlo time. To analyze the efficiency of this update, we choose $\Omega=0.6$ and $T=0.2$, as this choice of parameters belongs to a regime where the rod diffusion update has a high relative probability (Fig. 6(a)), and the other updates are not efficient (seen in acceptance probabilities in Figs. 3(b),4(b)). Our results are presented in Fig. 6(b), and we see that the local updates are not able to lose memory of the initial condition as the variance does not decay significantly, whereas using the rod diffusion update the variance drops to zero rapidly. The small residual value of the variance is due to the finite number of Markov chains which we have used (as $\braket{n_{i}}_{av}=\bar{n}$ only in the limit of averaging over an infinite number of Markov chains).
@@ 3.6 Benchmarking and autocorrelation functions
To ensure that we have implemented the detailed balance conditions correctly, we perform a comparison with exact diagonalization (ED) for a small system size. For the $L=2$ periodic lattice shown in Fig. 1, the number of dimer occupation states allowed by the Rydberg constraint are $2649$. This number is small enough to perform a full diagonalization, and thus to calculate energy at finite temperature. We carry out a comparison of the energy for a range of $\Omega$ as shown in Fig. 8(a), and find that the agreement between ED and QMC is of the order of $0.01\%$, and within the error bars of the QMC simulation. Thus we can conclude that the QMC simulations are unbiased and accurate. Note that in Fig. 8(a), we have used $T=0.006$, which is $\ll\Omega$ for the range of $\Omega$ studied.
@@ 4 Phase diagram of Kagomé link lattice
We now study the Kagomé link lattice using ED for the $L=2$ (24 sites) periodic lattice to gain an understanding of the phase diagram and study larger system sizes using the QMC updates discussed in the previous section. The latter allow us to extract the phase diagram in the thermodynamic limit. The Rydberg atom array on this lattice has been discussed in Ref. [1], where it was found that in the limit of $\Omega\gg\delta$, the ground state is a trivial paramagnet in the language of spin-1/2 particles, and that by lowering $\Omega$, one can cross into a quantum spin liquid phase and eventually into a valence bond solid phase with a large unit cell. These conclusions were reached by carrying out DMRG simulations on quasi-1D cylindrical geometries, and we would like to study the temperature dependence using our QMC directly in the 2D limit. The effect of temperature is most easily understood in the classical limit, where at high temperature, the system accesses all Rydberg allowed configurations, and at $\beta\gg\delta$, the system only accesses the manifold of states which have maximal density. This set is still macroscopically degenerate due to the inherent frustration of the underlying lattice. Once a quantum fluctuation is added, we expect a second energy scale controlled by both $\Omega$ and $\delta$, where the residual classical entropy is released, leading to a ground state with at most $O(1)$ degeneracy. In the phase diagram discussed below, we thus expect to find four phases. At high temperature, we expect a simple paramagnetic phase with maximal entropy and no signature in the order parameters we use to distinguish the other phases. At low temperature and $\Omega\gg\delta$, we expect the quantum paramagnet, which differs from the the high temperature one only in terms of entropy as it should have zero entropy for temperatures below the energy gap. At $\Omega\ll\delta$, we expect the classical spin liquid for $\Omega<T<\delta$, which has a non-zero entropy which is lesser than the maximal possible. In addition to the entropy, we also use a string order parameter to distinguish between the classical spin liquid and the quantum paramagnet as the former shows a non-zero value for strings of significant size whereas the latter does not. The quantum spin liquid is expected to emerge when we lower the temperature to $T\ll\Omega$ starting from the classical spin liquid regime, and this is characterized by zero entropy and a string order parameter behavior similar to that of the classical spin liquid. We do not find evidence of this last phase in the phase diagram which we are able to generate using the QMC simulations.
@@
Using ED, we study the specific heat as a function of both $\Omega$ and $\beta$ (for convenience we set $\delta=1$). We can trace the lower release in entropy as a peak in the specific heat, and we show this in Fig. 9(a). We find that this peak moves smoothly from $T\approx 10^{-5}$ at $\Omega=0.3$ to temperatures of $O(1)$ at high $\Omega$. This suggests that the curve traced by the location of the peak in Fig. 9(a) ends at $T=0$ for $\Omega=0$, which would be the natural expectation. At $\Omega\approx 1$, the distinction between the two peaks in specific heat which signify the release in entropy due to classical constraints and quantum fluctuations respectively, is lost. This is expected for $\Omega\approx\delta$ as the separation between scales required for two distinct peaks is lost. This phase diagram suggests that the quantum paramagnet is the correct ground state for all $\Omega>0$. This is in contrast with the expectation of a quantum spin liquid phase. We find that the low lying spectrum for small $\Omega$ has a unique ground state with a roughly equally spaced spectrum above it. This is also inconsistent with a $Z_{2}$ spin liquid on a torus, where one would expect four low lying states which have a topological nature, and are well separated from the rest of the spectrum.
@@
We show our results in Fig. 9(b) and (c), and find that these are similar to the $L=2$ lattice for system sizes of $L=4$ and $L=8$ down to temperatures of $\approx 0.01$. For low temperatures, calculating $C_{v}$ using the fluctuation- dissipation relationship, i.e. $C_{v}=\beta^{2}(\braket{E^{2}}-\braket{E}^{2})$, has large error bars, and thus we obtain $C_{v}$ directly by taking a numerical derivative of the energy. For our simulation ranges, this ensures an error bar of $<10\%$. The smallest value of $\Omega$ at which this temperature is low enough to detect the second peak in specific heat is $0.7$, and the temperature considered is 30 times smaller than the smallest scale in the system ($\Omega/2=0.35$). This point is consistent with the location of the transition to the spin liquid phase in Ref. [1]. If the spin liquid phase indeed appears close to this transition, it is not seen down to temperatures much smaller than $\Omega,\delta$.
@@ Appendix C Details of rod diffusion update
Each spatial site on the Kagomé link lattice can be visualized as a dimer between two neighboring sites on the Kagomé lattice, and it can be moved by pivoting it about one of its ends, as shown in Fig. 5(a). By extending this into the temporal direction, this pivot corresponds to a movement of a rod. The spatial site of the Kagomé lattice about which this pivot is carried out is chosen to not be in the membrane of modification already developed, as this would lead to the update undoing its own moves. This is similar to the spatial worm update [35], where dimers are pivoted around an exit site, which is not the site via which the worm entered the dimer. For the initial rod, this site is chosen at random, as there is no pre-existing membrane. An example of the membrane of modification is shown in Fig. 5(b). For the Kagomé lattice, a pivoting dimer has three possible options for its new position, as shown in Fig. 11. To decide which of these options to choose, we consider the corresponding rod configuration. We calculate the number of conflicting rods generated by moving the current rod to each of these three spatial positions while keeping the temporal ends of the rod the same. We choose the position with the least number of conflicting rods, and denote this number by $n_{c}$. If there are multiple such positions, we store the number of them by denoting them as the number of forward options ($p_{f}$), and choose one at random with uniform probability. If $n_{c}>1$, detailed balance would require a reverse proposed move which coordinates the movement of $n_{c}$ rods into the relevant positions, leading to the reverse movement of the current rod. This would imply that we would have to engineer an update which branches, and carries movements of multiple rods at the same step of the update. This would significantly complicate the update, and we cannot visualize a simple way to do this currently. An example of this sort of movement is shown in Fig. 5(c) and (d), where the movement of the current rod (shown in red) to its new position (shown in blue) creates a conflict with two rods. As shown, these two rods can be on the same spatial site or on different ones.
Figure 11: A single move within the rod diffusion update shown by projecting it onto the spatial plane. Circled are the numbers of rods at each spatial site which have a temporal overlap with the rod being moved (shown as gray dimer). This move can proceed as proposed as the number of legal forward options ($=1$), is the same as the number of legal backward options ($=1$).
@@
Thus if $n_{c}>1$, we abort the update. If not, we consider the reverse move, where the rod is being moved from its future position to its current position. In the dimer language, this is now the reverse pivot. Following the same procedure to pick the new position of the rod leads to a calculation of the number of backward options $p_{b}$. Detailed balance now requires that the probability of the forward process, $1/p_{f}$, be equal to that of the reverse process, $1/p_{b}$. This trivially translates to requiring $p_{f}=p_{b}$. If this condition is not met, we abort our update. We relax this condition only for the initial and final rod movement, where we require that $p_{f}^{start}=p_{b}^{end}$ and $p_{b}^{start}=p_{f}^{end}$. A sample pivot process which satisfies detailed balance is shown in Fig. 11. If the update is accepted, we update the relevant operators. With each rod movement, we carry all operators ($H_{K}$ and $H_{V}$ type) which are living on it to the next position inhabited by the rod. The update terminates when the current rod is in a position where there are no more conflicting rods to move. A short pseudo-code for this update is presented below.